Question 26 (1 point) Find the focus of the parabola y2 – 4y – 2x – 4 = 0. on
Question 27 (1 point) Find the exact value. CSC arccos
Question 28 (1 point) ã = (0.5, -3) and Ď = (-4,1). 112 à – 35 11. 4111 3v17 2013 5/10
Answer #1

Y4ya2-40 yny = 24+4 y-yyty = 204444 (4-282 = 26a44) comparing this egon to y-b2 = 2 planas 2022 spal , a=-4 and ba2 The focus8 9 = 20.5-37 and 3 = [~4,13 11 (25( 12- - (ا -رنل ال 96 - د ۱۱ -۱۱-13 -11) رہ جو \\= = 250